OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)?

OpenStudy (agent0smith):

Multiply them. (5x − 2)*log10 x

OpenStudy (anonymous):

the 10 after log is supposed to be down

OpenStudy (anonymous):

like logv10

OpenStudy (anonymous):

then x

OpenStudy (agent0smith):

I know

OpenStudy (agent0smith):

That makes no difference here though

OpenStudy (anonymous):

i still dont understand how to get to the answer from that

OpenStudy (btaylor):

Are you sure it isn't supposed to become f(g(x))? Because then it would be log10 (5x-2)

OpenStudy (btaylor):

Is it a closed or open circle?

OpenStudy (agent0smith):

if f(x) = log x and g(x) = 5x − 2 then f(x) • g(x) is just them two multiplied together. f(x) • g(x) = (5x − 2) log x that is your answer (you can distribute and simplify but it's fine like that imo). Unless as @BTaylor asked, you haven't posted it correctly

OpenStudy (btaylor):

That is correct and very clear. Thanks, @agent0smith.

OpenStudy (anonymous):

those arent in the answer choices theres f(x) • g(x) = log10 x^(5x − 2) f(x) • g(x) = log10 (5x − 2)^x f(x) • g(x) = 5x log10 x + 2 log10 x f(x) • g(x) = 2 log10 x − 5x log10 x

OpenStudy (agent0smith):

Then hopefully you know how to use some log laws... \[\Large a \log x = \log x^a\]

OpenStudy (agent0smith):

\[\Large (5x − 2) \log x = ...?\]

OpenStudy (agent0smith):

Hint.. the (5x-2) is just like the a in the previous formula.

OpenStudy (anonymous):

so would it be b?

OpenStudy (agent0smith):

Close, but look at the formula above again and you might see your mistake